# Quantifying the Fermi paradox via passive SETI: a general framework

### v1

@@ Abstract
In this paper we consider the extent to which a lack of observations from SETI may be used to quantify the Fermi paradox. Building on previous research, we construct a geometrical model to compute the probability of at least one detection of an extraterrestrial electromagnetic (EM) signal of galactic origin, as a function of the number $N$ of communicative civilizations. We first show that the probability of detecting a single signal is $\approx 0.6\delta/R$, where $\delta$ is the distance between the initial and final EM signals and $R$ is the radius of the Milky Way, for $\delta/R\ll 1$. Then we combine this analysis with the Drake equation $N=\mathcal{N}\delta/c$, where $c$ is the speed of light; this implies, applying a simplified toy model as an example, that the probability of detecting at least one signal is $>99\%$ for $\delta/c\gtrsim 10^{2.8}$ years, given that $\mathcal{N}=1$. Lastly, we list this toy model’s significant limitations, and suggest ways to ameliorate them in more realistic future models.
@@
SETI, Drake equation
pubyear: 2025pagerange: Quantifying the Fermi paradox via passive SETI: a general frameworkQuantifying the Fermi paradox via passive SETI: a general framework
@@ 1 Introduction
As motivation, we can begin with a general question. If there are currently $N$ signals from communicative civilizations in our galaxy, how likely it is that one would have been observed since the start of SETI? Suppose that we consider values of $N$ that seem reasonably conservative, but we find that the probability of having found a signal to-date from at least one of these sources is vanishingly small, given some simplifying assumptions; one may then argue that passive searches may not be meaningful. If this probability is large, however, one may argue that a lack of results from SETI is scientifically significant. In this paper, we will explain the general conditions under which latter scenario prevails. Further, we will explain how one can use such a result to constrict the parameter space of the Drake equation. This is in contrast to the study of the Drake equation in (Prantzos, 2020), in which the parameter space of the Drake equation is constrained via civilizations physically expanding their spheres of influence throughout the galaxy.
@@
The reader may wonder what we mean by “some simplifying assumptions”. This paper will define and discuss several such assumptions, and we will leave it to the reader to decide their plausibility. We hope that this paper will provide a framework in which to consider interpretations and resolutions of the Fermi paradox, within the context of past, present, and future searches.
We will first discuss the geometrical interpretation of this problem. Then, we compute the probability of a single detection via this model, followed by a discussion of how to compute the probability of at least one observation. This discussion will build upon Grimaldi (2017); we provide an exact solution in the two-dimensional case and a derivation of a useful approximation for the probability of one observation, and we show that in the simplest interpretation the location of Earth does not affect the likelihood of observation. The latter provides a very useful way of generalizing these results to more complex probability distributions. Our treatment will differ further in that we then connect these results to the Drake equation, and argue that its parameter space can be constrained via a null result from SETI. This is the crux of our work. In this paper, we will refer to “communicative civilizations” as “civilizations”.
2 A Simple Two-Dimensional Geometrical Model
2.1 Introduction
These light fronts create an annulus whose area may only be partly located within the MW galaxy. Let us discuss how to compute this. The location of the source is $(r_{0},\theta_{0})$, and Earth is denoted by the blue dot, of position $(r_{E},\theta_{E})$. Observation will correspond to an orientation where the Earth is in-between the two light-fronts. Hence, observation will only occur when
Figure 1: Here, we depict the basic geometry of our model. The MW galaxy has a radius “R”. The two light-fronts, emanating from an extraterrestrial source “S”, are labelled “A” and “B”. The source is a distance $r_{0}$ from the center of the MW galaxy, and $\tilde{r}_{A}$ is the distance from the origin to a point on light-front A. The distance of the Earth from the center of the Milky Way is labeled “$r_{E}$”, and this figure depicts the condition defined by Equation 1.
For a single pair of light fronts, the probability of observation for a randomly-placed observer is the probability that this randomly-placed observer will fall in-between the light fronts; this is $\tilde{A}/\pi R^{2}$, where $\tilde{A}$ is the area that is both in-between the two light fronts and also within the MW galaxy. The area $\tilde{A}$ is circumscribed by the curves whose intersections with the edge of the MW galaxy are defined by the red and green dots. All points along the light-front $A$ ($B$) form angles $\tilde{\theta}_{A}$ ($\tilde{\theta}_{B}$), as shown in Figure 1. The distance from the center of the MW galaxy to any point along light-front $A$ ($B$) is $\tilde{r}_{A}$ ($\tilde{r}_{B}$). The red (green) dots are defined by $\tilde{r}_{A}=R$ ($\tilde{r}_{B}=R$), which implies that (keeping to just the red dots as an example)
$r_{A}^{2}=R^{2}+r_{0}^{2}-2Rr_{0}\cos(\theta_{0}-\tilde{\theta}_{A}),$ (2)
and thus
We can split useful $(r_{0},r_{A})$ parameter space into two parts: (1) where the argument of $\arccos$ is within $[-1,1]$, and (2) where it is not. In the first case we will have two [^1] Except the singular case where $\Delta\theta=0$ when the argument is $1$. values for $\tilde{\theta}_{A}$, which we will call $\tilde{\theta}_{A1}=\theta_{0}-\Delta\theta$ and $\tilde{\theta}_{A2}=\theta_{0}+\Delta\theta$. This implies that
$R+r_{0}>r_{A}>R-r_{0}.$ (4)
2.2 Annulus area and probability of observation
In Figure 2, we depict just the essential components of our geometrical model sufficient to compute $\tilde{A}$. The inner area, $\tilde{A}_{A}$, can be bifurcated into two parts: the area along the inner concentric circle from one red dot to the other (shaded with red triangles), and the area along the edge of the MW (shaded with green x marks) from one red dot to the other. The latter is $R^{2}\left(\tilde{\theta}_{A2}-\tilde{\theta}_{A1}\right)/2=R^{2}\Delta\theta=\Delta\theta$, taking $R=1$; the former is $r_{A}^{2}\psi$, where $\psi$ is (as depicted in Figures 1 and 2) the angle between $r_{0}$ and $r_{A}$; $\psi$ appears to be
$\frac{\sin\psi}{\tilde{r}_{A}}$ $=\frac{\sin\Delta\theta}{r_{A}},$
$\Rightarrow\psi$ $=\sin^{-1}\left(\frac{\sqrt{4r_{0}^{2}-\left(1+r_{0}^{2}-r_{A}^{2}\right)^{2}}}{2r_{0}r_{A}}\right)\equiv\sin^{-1}\left(\chi\right)$
where we have taken $\tilde{r}_{A}=1$ so that $\psi$ is defined as in Figure 2. This angle isn’t quite right, however, since $\chi$ approaches 00 as $r_{A}$ approaches $1-r_{0}$. As we can see from Figure 2, however, $\psi$ should be $\pi$ in this limit. Thus,
$\psi=\gamma\pi+\alpha\sin^{-1}\left(\chi\right),$ (5)
where $\alpha=-1$ ($\gamma=1$) when $\psi>\pi/2$ and $+1$ (00) otherwise. When these two regions meet at $\psi=\pi/2$,
$\sin^{-1}\left(\chi\right)=\pi-\sin^{-1}\left(\chi\right),$
which implies that $\chi=1$ and
$r_{A}^{2}+r_{0}^{2}+1-2r_{A}-2r_{0}+2r_{A}r_{0}=0.$
Equation 2 implies that $r_{A}=\sqrt{1+r_{0}^{2}-2r_{0}\cos\Delta\theta}$, which, when input into the equation just above, gives $r_{0}=\cos\Delta\theta$. Thus, $r_{A}=\sqrt{1+r_{0}^{2}-2r_{0}^{2}}=\sqrt{1-r_{0}^{2}}$ when [^2] We can see this also via the Pythagorean theorem, since the triangle in Figure 2 is a right triangle when $\psi=\pi/2$. $\psi=\pi/2$. Thus, $\alpha=-1$ ($\gamma=1$) when $r_{A}<\sqrt{1-r_{0}^{2}}$ and $+1$ (00) otherwise.
$\tilde{A}_{A}=\Delta\theta+r_{A}^{2}\psi-r_{0}\sin\Delta\theta.$ (6)
Therefore, the probability of a single observation is
$P=\begin{cases}\left(\tilde{A}_{B}-\tilde{A}_{A}\right)/\pi,\hskip 14.22636pt1+r_{0}>r_{A}>1-r_{0},\\ r_{B}^{2}-r_{A}^{2},\hskip 36.98866ptr_{A}\leq 1-r_{0},\\ \end{cases}$ (7)
Figure 2: Here, we depict the three areas that compose $\tilde{A}$: the area denoted by red triangles ($\psi r_{A}^{2}$), the area denoted by green x symbols ($\Delta\theta$), and their overlap. Thus, we subtract the last area from the sum of the first two.
For the rest of this paper, we will explore the mechanics of this model via a toy model. This model is based upon several approximations, and, thus, the exact results are superfluous. Since for many practical purposes $r_{B}-r_{A}$ will be small compared to the radius of the MW, for example, we will in the next section implement this approximation and use it throughout the paper.
2.3 $\delta\equiv r_{B}-r_{A}\ll 1$ regime
Let us now approximate Equation 7 where $\delta\ll 1$. We begin with the bifurcation of the parameter space into whole source circles and partial source circles, as described in section 2.1. When $\delta\equiv r_{B}-r_{A}\ll 1$, we may approximate the area $\tilde{A}$ as $\tilde{A}\approx\Omega\cdot\delta r_{A}$, where $\Omega=2\pi$ when the annulus is completely within the MW and $\Omega=2\psi$ otherwise. (See Equation 5 for the definition of $\psi$.) Thus,
$+2\pi\cdot\zeta[-rA+1-r0]\bigg{)},$ (8)
where
$\zeta[x]=\begin{cases}1,\hskip 28.45274ptx>0,\\ 0,\hskip 28.45274ptx<0.\\ \end{cases}$ (9)
The expression $\zeta[rA-\left(1-r0\right)]\zeta[-rA+\left(1+r0\right)]=\zeta_{p}$ enforces the first conditional in Equation 7 (i.e., partial circles), and $\zeta[-rA+1-r0]=\zeta_{w}$ enforces the second in Equation 7 (i.e., whole circles). The average area over the MW is, therefore,
[…]
×$\times\int_{0}^{2}\int_{0}^{1}r_{0}A_{\delta}dr_{0}dr_{A}$
$=0.6\pi\delta,$
$\Rightarrow P\approx P_{\delta}$ $=0.6\delta,$ (10)
2.4 The effect of the observer’s location
which is satisfied for all $r_{E}$. We can go further and use this to recompute Equation 2.3. The probability of observation is the likelihood that an observer will fall in-between the light fronts $r_{A}$ and $r_{A}+\delta$, and, since the domain of $r_{A}$ is $[0,1+r_{0}]$, this probability is $\delta/\left(1+r_{0}\right)$ if we assume that all $r_{A}$ are equally likely. Defining $\zeta=\zeta[-rA+\left(1+r_{0}\right)]$, the average probability for small $\delta$ is therefore
[…]
$=0.6\delta.$
In the next section, we discuss how to compute the probability of at least one observation, as a function of the probability of observing $N$ probabilities $P_{1}$, $P_{2}$, …,$P_{N}$. Since we seek in this paper to explain the basic mechanics of the model, we will just take $P=0.6\delta$ throughout the remainder of this work.
3 Probability Distributions
3.1 Introduction
In this introductory section, we follow (Grimaldi, 2017). For $N$ civilizations having independent detection probabilities $P_{1}$, $P_{2}$, …,$P_{N}$, the probability of no detection is
Thus, the probability of at least one detection is
$\mathcal{P}=1-\prod_{i=1}^{N}\left(1-P_{i}\right).$ (12)
One may analyze the probability distributions $P_{i}$ as in Kreifeldt (1971) and Wallenhorst (1981). Since our goal is to provide a highly general introduction to our model, however, in this paper we will consider the greatly simplified case where each probability $P_{i}=P$ is identical. We refer to this as the homogeneous probability model, which we can interpret as a toy model.
3.2 The homogeneous probability model
$\mathcal{P}=1-\left(1-P\right)^{N}.$
Solving for $\log_{10}N$, we obtain
3.2.1 Approximating $\mathcal{P}$ in the $\delta\ll 1$ regime.
In the $\delta\ll 1$ regime, the probability of a single observation $P$ will be $\ll 1$. This conclusion follows from Section 2.3. We can now approximate Equation 14 in this regime, starting from Equation 13. This yields
where in the first line we have assumed that $P\ll 1$. This result is superposed on the exact result (Equation 14) in Figure 3, in which it is depicted as green dashed lines for $\mathcal{P}=0.1$, $0.5$, and $0.99$. Equation 14 is plotted with via the solid curves. From this figure, we can see that Equation 3.2.1 is accurate to at least an order of magnitude so long as $\delta\lesssim 1$.
Figure 3: In Figure 3, we depict $\log_{10}N$ vs. $\log_{10}P$ (Equation 14) in the homogeneous probability model, for $\mathcal{P}=0.1$, $0.5$, and $0.99$. These curves are overlayed with the small-$P$ approximation Equation 3.2.1, shown in green dashed lines, for each $\mathcal{P}$.
4 Numerical Analysis
In Table 4, we publish the probability of a single observation, computed using three different methodologies: $P_{exact}$, via Equation 7; $P_{approx}$, via Equation 2.3; and, $P_{test}$, via Equation 1. The first is computed by splitting the domains of $r_{0}$ and $r_{A}$ into $10^{3}$ segments each for $\delta=10^{-1},10^{-1}$, and $10^{-3}$, and $10^{4}$ segments each for $\delta=10^{-4}$. The second is trivially derived. The third is calculated by splitting the domains of $r_{0}$, $r_{A}$, and $\theta_{0}$ into $10^{3}$ segments each for $\delta=10^{-1},10^{-1}$, and $10^{-3}$, and $10^{4}$ segments for $\delta=10^{-4}$. This also requires one to choose $r_{E}$ and $\theta_{E}$ values; these, however, do not affect $P_{test}$. See Section 2.4 for a discussion.
Table 1: Below we list observation probabilities computed three different ways. One is derived from Equation 1, computed by iterating $r_{0}$, $r_{A}$, and $\theta_{0}$ values over their domains. We refer to this as $P_{test}$. Probexact is computed by inserting iterated $r_{0}$ and $r_{A}$ values into Equation 7, while Probapprox is computed trivially via Equation 2.3.
[…]
$10^{-1}$ $0.632526$ $0.600000$ $5.1$ $0.636311$
$10^{-2}$ $0.603640$ $0.600000$ $5.8\times 10^{-1}$ $0.603781$
$10^{-3}$ $0.600549$ $0.600000$ $8.7\times 10^{-2}$ $0.600724$
$10^{-4}$ $0.600055$ $0.600000$ $3.4\times 10^{-2}$ $0.599790$
The percent error in Table 4 refers to the error of $P_{approx}$ with respect to $P_{exact}$. We now turn to the meaning of our model in the context of the Fermi paradox and SETI.
5 Discussion
In this section, we will apply the geometrical model described in the previous section and discuss its implications. First, we will describe how the Fermi paradox can be quantified via SETI, and then we will discuss the limitations of the toy model we present here.
5.1 The Fermi paradox in light of SETI
The model we’ve articulated here can be used to quantify the Fermi paradox, using the fact that SETI has not observed an EM signal. Its usefulness is best understood via the Drake equation (Drake (1965)), which is
@@ 5.1 The Fermi paradox in light of SETI
Since $P\approx 0.6\delta$ for small $\delta$, we can write Equation 3.2.1 as
@@
We can now write $\delta=r_{A}-r_{B}=lc/R$, where $c$ is the speed of light and $R$ is the radius of the MW galaxy in light-years. We will use years and light-year interchangeably, thus taking $c=1$, and we take $R=5\times 10^{4}$. Thus, $\delta=\frac{1}{5}\cdot 10^{-4}l$. This gives
@@
$\approx-\log_{10}l+\log_{10}\left(\frac{25}{3}\cdot 10^{4}\cdot\ln\left(1-\mathcal{P}\right)^{-1}\right)\equiv-\log_{10}l+\mathcal{C}(\mathcal{P}).$
@@
Let us now presume that $\mathcal{N}\lesssim 1$, as argued in the original formulation of the Drake equation Drake (1965). This precludes the area above the red line in Figure 4. The area to the top right of the blue line in that figure implies $\mathcal{P}>0.99$. Suppose we restrict ourselves to SETI results which have $>1\%$ probabilities of occurring; this precludes the aforementioned area. In the original iteration, Drake considered $l=10^{3}$ to $10^{4}$ years as reasonably conservative values. We can see from Figure 4, however, that, if each civilization emits EM signals for $l\approx 10^{2.8}$ years, there is about a $99\%$ probability that we would have observed at least one of these signals, given $\mathcal{N}\approx 1$. The latter presumption reduces Equation 5.1 to
@@
Figure 5: Here, we plot $\log_{10}l$ vs. $\mathcal{P}$, as per Equation 18. Thus, here we presume that $\mathcal{N}\approx 1$.
These results, under the limitations described below, imply that it is unlikely that there are more than $\mathcal{O}(10^{3})$ communicative civilizations unless $\mathcal{N}$ is appreciably smaller than $1$. This is depicted in Figure 5. Since $\mathcal{N}$ is multivariate, this constraint cannot in and of itself restrict any of the individual parameters which define $\mathcal{N}$. There are, however, some independent constraints on some of those aforementioned parameters. For instance, $r_{*}$ is likely (Elia et al. (2022)) around $1$, and it currently seems likely that $f_{p}$ is (Cassan et al. (2012)) around $1/2$. If these hold, then $n_{e}f_{l}f_{i}f_{c}$ should likely be at least around an order of magnitude less than $1$ if $l\gtrsim\mathcal{O}(10^{3})$ and $\mathcal{N}\approx 1$. The model we have discussed here has rather notable limitations, and so this argument is far from dispositive. Perhaps with significantly more data and a more realistic model of this general type, however, more definitive statements along these lines might be made.
We now turn to the limitations of this model, and suggestions for how to improve it.

### v2

# The Open Journal of Astrophysics

@@ Abstract
In this paper we consider the extent to which a lack of observations from SETI may be used to quantify the Fermi paradox. We compute the probability of at least one detection of an extraterrestrial electromagnetic (EM) signal of Galactic origin, as a function of the number $N$ of communicative civilizations. We show how this is derivable from the probability of detecting a single signal; the latter is $\approx 0.6\delta/R$, where $\delta$ is the distance between the initial and final EM signals and $R$ is the radius of the Milky Way, for $\delta/R\ll 1$. We show how to combine this analysis with the Drake equation $N=\mathcal{N}\delta/c$, where $c$ is the speed of light; this implies, applying a simplified toy model as an example, that the probability of detecting at least one signal is $99\%$ for $\delta/c\gtrsim 10^{2.8}$ years, given that $\mathcal{N}=1$. Lastly, we list this toy model’s significant limitations, and suggest ways to ameliorate them in more realistic future models.
@@
SETI, Fermi paradox
@@ 1 Introduction
As motivation, we can begin with a general question. If there are currently $N$ signals from communicative civilizations in our galaxy, how likely is it that one would have been observed since the start of SETI? Suppose that we consider values of $N$ that seem reasonably conservative, but we find that the probability of having found a signal to date from at least one of these sources is vanishingly small, given some simplifying assumptions; one may then argue that passive searches may not be meaningful. If this probability is large, however, one may argue that a lack of results from SETI is scientifically significant. In this work, we will explain the general conditions under which the latter scenario prevails. Further, we will explain how one can use such a result to constrict the parameter space of the Drake equation. This is in contrast to the study of the Drake equation in (Prantzos, 2020), in which the parameter space of the Drake equation is constrained via civilizations physically expanding their spheres of influence throughout the galaxy.
@@
The reader may wonder what we mean by “some simplifying assumptions”. We define and discuss several such assumptions, and we will leave it to the reader to decide their plausibility. We hope that this paper will provide a framework in which to consider interpretations and resolutions of the Fermi paradox, within the context of past, present, and future searches. Here, we will refer to “communicative civilizations” as “civilizations”.
In the next section, we discuss the conceptual crux of this paper. We then conceptually introduce our geometrical model of the MW galaxy. This is followed by a discussion of how to compute the probability of at least one observation from the probability of individual observations, and a discussion of how one can use these results to constrain the Drake equation. Lastly, we interpret our work within the context of previous SETI and Fermi paradox research, and provide suggestions for fixing the limitations of the simple model described here.
1.1 Constraining the Drake Equation via SETI
The essential idea we wish to present is that a null SETI result can be used to constrain the Drake equation (Drake (1965))
@@ 1.1 Constraining the Drake Equation via SETI
The likelihood $P_{i}$ of observing a civilization’s signals depends on the number of years during which the civilization has emitted signals. We can then determine the probability $\mathcal{P}$ of observing at least one signal by imagining many civilizations. Therefore, the fact that SETI has not observed any life-produced signals can be used to place constraints on $l$. By combining an upper bound on $l$ with Equation 1, we can exclude parts of $\mathcal{N}-l$ parameter space, and, in this way, a null SETI result quantifies the Fermi paradox.
So as not to misguide the reader by focusing on mathematical details, we will give a brief and conceptual introduction to our geometrical model in the next subsection, sufficient to follow the connection between $P_{i}$ and the Drake equation. The details are provided in Appendix A.
1.2 Conceptual introduction to the geometrical model
In Figure 1, the basic features of our geometrical model are depicted. We model the MW as a two-dimensional disk of radius $R$, containing civilizations (labeled “SS”) located at distances $r_{0}$ from the center of the MW. The first and last signals have radii $r_{B}$ and $r_{A}$, respectively. The probability of observation of a single signal ($P_{i}$) is proportional to the signal area, denoted by yellow diamonds in the aforementioned figure. This area is derived in Appendix A.
For some applications, the exact area is not needed; in such applications, $r_{B}-r_{A}\equiv R\delta\ll R$ [^1] From here on, we will normalize $\delta$ as $\left(r_{B}-r_{A}\right)/R$ and take $R=1$, redefining $R$ in units of distance only as necessary.. Since our primary goal in this paper is to conceptually describe how SETI can be used to constrain the Drake equation, we will for simplicity take $\delta\ll 1$ throughout. In Appendix A, we show that in this regime the probability of observing a signal is $P_{i}=0.6\delta$.
In Table 2, we list the primary parameters used in the paper, summarize their meanings, and direct the reader to the equation or section in which they are defined. Note that there are three ways in which we will compute $P_{i}$. The aforementioned approximation for $P_{i}$ when $\delta\ll 1$ we call $P_{\delta}$. Contrarily, the exact probability (Equation 14) we refer to as $P_{exact}$. Finally, $P_{test}$ is calculated by iteratively generating sources, and explicitly determining if each source is observable. This is discussed in Section 3. Next, we will discuss how to connect $N$, $P_{i}$, and $\mathcal{P}$.
Figure 1: This figure conceptually describes the geometrical model used in this paper. The essential idea is that the greater the number of years during which a civilization has emitted signals, the greater the likelihood of observing those signals. This likelihood is proportional to the signal area, which is denoted with yellow diamonds. Here, the signal is being sent from a source “SS”, which is a distance $r_{0}$” from the center of the MW. The first and last signals have radii $r_{B}$ and $r_{A}$. Thus, $l$ in Equation 1 can be connected to the probability $\mathcal{P}$ of observing at least one signal.
Table 1: Here we summarize all of the primary symbols we use in this paper, and provide links to the sections in which they are defined.
Symbol Meaning See
$\vskip 3.41418ptl$ mean lifetime of civilizations (years), $r_{B}-r_{A}$, where $c=1$, for a single civilization Section 4.1
$\vskip 1.13791pt\delta$ $\vskip 1.13791ptl/R$ Section 1.2
$\vskip 2.27626ptP_{exact}$ the exact probability of a single observation Equation 14
$\vskip 1.42271ptP_{\delta}$ $\vskip 1.13791pt0.6\delta$ Equation A.3
$\vskip 8.5359ptP_{test}$ the probability of a single observation, via iterating through all EM sources and directly checking the observation condition Equation 8
$\vskip 1.42271pt\mathcal{P}$ the probability of at least one SETI observation Section 2.1
$\vskip 1.42271pt\mathcal{N}$ all factors in the Drake equation except $l$ Sections 4.1 and 1
2 Connecting SETI to our Geometrical Model
In this section, we will discuss how to connect a null SETI result to $P_{i}$ and $\mathcal{P}$. We will then simplify our discussion by taking $\delta\ll 1$, and by considering only a homogeneous distribution of $P_{i}$.
2.1 Probability distributions
In this subsection, we discuss how to determine the probability of at least one detection $\mathcal{P}$ as a function of the distribution $P_{i}$. This is similar to the discussion in (Grimaldi, 2017). For $N$ civilizations having independent detection probabilities $P_{1}$, $P_{2}$, $P_{N}$, the probability of no detection is
and, thus, the probability of at least one detection is
$\mathcal{P}=1-\prod_{i=1}^{N}\left(1-P_{i}\right).$ (2)
One may analyze the probability distributions $P_{i}$ as in Kreifeldt (1971) and Wallenhorst (1981). Since our goal is to provide a highly general introduction to our model, however, we take the greatly simplified case where each probability $P_{i}=P$ is identical. We refer to this as the homogeneous probability model, which we can interpret as a toy model.
2.2 The homogeneous probability model
$\mathcal{P}=1-\left(1-P\right)^{N}.$
Solving for $\log_{10}N$, we obtain
2.3 Approximating $\mathcal{P}$ in the $\delta\ll 1$ regime.
In the $\delta\ll 1$ regime, the probability of a single observation $P$ will be $\ll 1$. This conclusion follows from Section A.3. We can now approximate Equation 4 in this regime, starting from Equation 3. This yields
where in the first line we have assumed that $P\ll 1$. This result is superposed on the exact result (Equation 4) in Figure 2, in which it is depicted as green dashed lines for $\mathcal{P}=0.1$, $0.5$, and $0.99$. Equation 4 is plotted via the solid curves in red, blue, and black for the same $\mathcal{P}$ values. From this figure, we can see that Equation 2.3 is accurate to at least an order of magnitude so long as $\delta\lesssim 1$.
Figure 2: Here, we depict $\log_{10}N$ vs. $\log_{10}P$ (Equation 4) in the homogeneous probability model, for $\mathcal{P}=0.1$ (in red), $0.5$ (in blue), and $0.99$ (in black). These curves are overlayed with the small-$P$ approximation Equation 2.3, shown in green dashed lines, for each $\mathcal{P}$.
3 Numerical Analysis of Our Geometrical Model
[…]
Table 2: Below we list observation probabilities computed three different ways. $P_{test}$ is derived from Equation 8, computed by iterating $r_{0}$, $r_{A}$, and $\theta_{0}$ values over their domains and checking to see if the observation condition is met. See Figure 5 for the definitions of these parameters. Pexact is computed by inserting iterated $r_{0}$ and $r_{A}$ values into Equation 14, while Pδ is computed trivially via Equation A.3. The percent error in Table 2 refers to the error of $P_{\delta}$ with respect to $P_{exact}$.
$\delta$ $P_{exact}$ $P_{\delta}$ $\%$ error $P_{test}$
$10^{-1}$ $0.632526$ $0.600000$ $5.1$ $0.636311$
$10^{-2}$ $0.603640$ $0.600000$ $5.8\times 10^{-1}$ $0.603781$
$10^{-3}$ $0.600549$ $0.600000$ $8.7\times 10^{-2}$ $0.600724$
$10^{-4}$ $0.600055$ $0.600000$ $3.4\times 10^{-2}$ $0.599790$
We now turn to the meaning of our model in the context of the Fermi paradox and SETI.
4 Discussion
In this section, we will apply our geometrical model to the Drake equation and discuss its implications. Finally, we will discuss the limitations of the toy model we present here.
4.1 The Fermi paradox in light of SETI
Since $P\approx 0.6\delta$ for small $\delta$, we can write Equation 2.3 as
@@ 4.1 The Fermi paradox in light of SETI
We can now write $\delta=lc/R$, where $c$ is the speed of light and $R$ is the radius of the MW galaxy in light-years. We will use years and light-year interchangeably, thus taking $c=1$, and we take $R=5\times 10^{4}$. Thus [^4] Recall that in Section 1.2 we took $c=1$ and $R\delta=r_{B}-r_{A}$; therefore, $\delta$ is the normalized distance that the EM signals have travelled through the MW galaxy, consistent with our result here., $\delta=\frac{1}{5}\cdot 10^{-4}l$. This gives
@@
$\approx-\log_{10}l+\log_{10}\left(\frac{25}{3}\cdot 10^{4}\cdot\ln\left(1-\mathcal{P}\right)^{-1}\right)$
$\equiv-\log_{10}l+\mathcal{C}(\mathcal{P}).$
@@
Let us now presume that $\mathcal{N}\lesssim 1$, as argued in the original formulation of the Drake equation Drake (1965). This precludes the area above the red line in Figure 3, in which we plot Equation 4.1. The magenta line represents $\mathcal{N}=0.1$, added for readability. The area to the top right of the blue line in that figure implies $\mathcal{P}>0.99$, since the blue line represents $\log_{10}\mathcal{N}$ vs $\log_{10}l$ for $\mathcal{P}=0.99$. Suppose we restrict ourselves to SETI results which have $1\%$ probabilities of occurring; this precludes the aforementioned area. In the original iteration, Drake considered $l=10^{3}$ to $10^{4}$ years as reasonably conservative values. We can see from Figure 3, however, that, if each civilization emits EM signals for $l\approx 10^{2.8}$ years, there is about a $99\%$ probability that we would have observed at least one of these signals, given $\mathcal{N}\approx 1$. The latter presumption reduces Equation 4.1 to
@@
Figure 4: Here, we plot $\log_{10}l$ vs. $\mathcal{P}$, as per Equation 7. Thus, here we presume that $\mathcal{N}\approx 1$. This shows that, given the limitations of our toy model, it is unlikely that there are about $10^{2.8}$ civilizations or more unless $\mathcal{N}$ is considerably smaller than $1$.
These results, under the limitations described below, imply that it is unlikely that there are more than $\mathcal{O}(10^{3})$ communicative civilizations unless $\mathcal{N}$ is appreciably smaller than $1$. This is depicted in Figure 4. Since $\mathcal{N}$ is multivariate, this constraint cannot in and of itself restrict any of the individual parameters which define $\mathcal{N}$. There are, however, some independent constraints on some of those aforementioned parameters. For instance, $r_{*}$ is likely (Elia et al. (2022)) around $1$, and it currently seems likely that $f_{p}$ is (Cassan et al. (2012)) around $1/2$. If these hold, then $n_{e}f_{l}f_{i}f_{c}$ should likely be at least around an order of magnitude less than $1$ if $l\gtrsim\mathcal{O}(10^{3})$ and $\mathcal{N}\approx 1$. The model we have discussed here has rather notable limitations, and so this argument is far from dispositive. Perhaps with significantly more data and a more realistic model of this general type, however, more definitive statements along these lines might be made. Let us turn toward the meaning of what we’ve done in relation to similar research.
4.2 Comparing our work to previous SETI and Fermi paradox research
Thus far we have not placed our work within the context of previous SETI and Fermi paradox research. What makes this paper unique is that we present a method by which the null SETI result can be used to impose constraints upon $l$; and, hence, we can constrain the parameter space of the Drake equation. Previous research (Forgan (2009); Maccone (2012); Miller & Felton (2017), Maccone (2012)) has presented numerical methods for simulating the formation of life in the MW galaxy, via Monte Carlo techniques and Bayesian analysis. Similar work has been done within particular Fermi paradox frameworks–for instance, within the context of the Great Filter hypothesis (Aldous (2012)). Statistical analysis has been used to argue that the Fermi paradox is not likely to be paradoxical at all (Sandberg et al. (2018)). Others (Westby & Conselice (2020)) have investigated the impact of astrophysical scenarios on the Drake equation, constraining its parameters via astronomical observations.
In this work, we aim to fill a rather conspicuous gap in the literature, by showing that SETI can be a useful tool by which to analyze the prevalence of civilizations within our galaxy. Merely by searching for life-produced signals and failing to find any, we can provide a quantitative interpretation of the Drake equation that, when combined with independent evidence, may one day provide humanity with a scientific understanding of its place within the Galaxy. The model we have presented here is intended, however, to give an example of how one might accomplish this; we are thus not naive about its limitations. In this next section, we turn to the limitations of this model and suggestions for how to improve it.